OpenStudy (anonymous):

FAN AND MEDAL! EASY QUESTION! is the 4th square root the same as to the 1/4 power? meaning that 2^1/2 = the forth squair root or 2?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@mathslover @phi @Callisto please help! its really just a technical question.

OpenStudy (anonymous):

@terenzreignz

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

2^1/2 = sqrt(2) 2^1/4 = 4rt(2) k^1/n = nrt(k)

OpenStudy (anonymous):

that's what I thought, thank you :)

OpenStudy (phi):

Here is how you know First, remember when you multiply numbers with the same base, you add their exponents. Example: \[2^1 \cdot 2^1 = 2^{1+1}= 2^2 \text{ or } 4 \] Second, by definition, the fourth root of 2 (call it a) means \[ a \cdot a \cdot a \cdot a= 2\] Third, let's multiply 2^(¼) times itself 4 times: \[ 2^{\frac{1}{4}} \cdot 2^{\frac{1}{4}} \cdot 2^{\frac{1}{4}} \cdot 2^{\frac{1}{4}} = 2^{\frac{1}{4}+\frac{1}{4}+\frac{1}{4}+\frac{1}{4}}= 2^{\frac{4}{4}}= 2^1=2 \] in other words, 2^(¼) must be the fourth root of 2, because when we multiply it by itself 4 times, we get 2